OpenStudy (this):

helpp plzzz what is the mean absolute deviation of the set (12,10,14,4,5)? a 18.0 b9.0 c 3.6 d 1.8

OpenStudy (anonymous):

put the data in order first

OpenStudy (anonymous):

4,5,10,12,14

OpenStudy (anonymous):

average,\[(4+5+10+12+14)\div5=9\] \[(\left| 4-9 \right|+\left| 5-9 \right|+\left| 10-9 \right|+\left| 12-9 \right|+\left| 14-9 \right|)\div5\]\[(5+4+1+3+5)\div5=3.6\] Thus,C

OpenStudy (anonymous):

Syderitic good job

OpenStudy (this):

@syderitic thanks a lot for ur nice help